# Understanding Iterative Combinatorial Auction Designs via Multi-Agent Reinforcement Learning

### v1

@@ 2.2. Finding Equilibria
At a high level, MARL algorithms vary along two key dimensions. The first is whether they use a lookup table or function approximation to represent players’ policies. In tabular methods, each information state in the game is represented separately; bidders faced with a new information state do not reason about its similarity to other states with which they have previous experience, but instead start learning afresh. This is a double-edged sword: it does not constrain players’ behavior in any particular way, allowing them to learn flexible strategies, but causes erratic behavior (typically close to uniform random) in very rarely encountered states. With function approximation (i.e., deep reinforcement learning), each information state is represented by a feature vector and the agent’s behavior is regularized to be similar in states having similar feature vectors. This allows agents to generalize from one part of the game tree to another, but can be more difficult to train reliably. Some tabular methods, such as Counterfactual Regret Minimization (Zinkevich et al., 2007), also assume perfect recall—that an agent never forgets anything that they previously observed—while function approximation methods generally do not. Consider whether it seems strategically important that agents never blend observations together.
The second key dimension distinguishing MARL algorithms is which parts of the game tree are explored in each training iteration. Many algorithms from single-agent reinforcement learning, such as Q-learning (Watkins and Dayan, 1992) or policy-gradient methods (Sutton et al., 1999), only sample a single path through the game tree at each iteration. While it is possible to use these algorithms in multi-agent settings by running them independently for each player, they can struggle due to focusing most of their training effort on a small fraction of the game tree. At the other extreme, algorithms such as Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007) enumerate every information state in each iteration, getting good coverage over the entire game tree, but their runtime quickly becomes infeasible for larger games. Lastly, some algorithms fit between these two extremes, enumerating one player’s actions but sampling opponent actions and chance outcomes; one such example is external-sampling Monte-Carlo CFR (MCCFR) (Lanctot et al., 2009).
If it is feasible to enumerate the entire game tree, we see little reason not to use a tabular method that explores much of the game tree, such as CFR or related variants; at the very least, we recommend such algorithms as a baseline for comparing to other approaches. For very large games, where it becomes infeasible to run even a single iteration of a CFR variant, function approximation methods are likely the way forward.
@@
Complex auctions typically admit multiple equilibria. When using non-deterministic MARL algorithms, a simple way to find multiple equilibria is to run the algorithm with multiple random seeds, which influence both its random initialization and any random events within the RL environment, causing it to take a different learning trajectory on each run. We thus recommend non-deterministic algorithms, e.g., preferring Monte-Carlo sampling variants of CFR over the original algorithm. We have less experience with deterministic MARL algorithms; we speculate about potential approaches to identifying multiple equilibria with such algorithms in Section 5.2.1.
When two or more actions produce the same terminal reward, RL algorithms have no reason to prefer one action over another, and equilibria multiply. This can be detrimental to both convergence and auction outcomes. For example, consider a bidder facing an incredibly strong opponent that will price them out of the auction no matter how they behave in early rounds. Since they cannot impact their own utility, they can play any strategy in equilibrium, ranging from immediately dropping out of the auction to raising their opponent’s prices up to their own value; their arbitrary choice will impact many auction outcomes of interest (e.g., auction length, revenue). We therefore advocate for adding small, “secondary” rewards to the RL environment, biasing agents toward certain strategies (e.g., preferring to drop out than to stay in) when all else is equal.
@@
An analyst must decide whether to focus solely on pure equilibria or also to consider mixed equilibria. There are several advantages to studying pure equilibria. First, working with pure equilibria is computationally beneficial: restricting to pure strategies greatly reduces the strategy space, tending to make it easier to train, validate, and evaluate policies. Second, some MARL algorithms may not be able to converge to mixed equilibria at all. For a broad class of MARL algorithms, the only stable points of the learning dynamics are pure equilibria (Flokas et al., 2020); we believe that a similar result is likely to hold for many algorithms, such as CFR. While there is no guarantee that every game will have any pure equilibria, in our own experiments, we have found that MARL algorithms can find a pure equilibrium in the vast majority of our games.
We recommend considering whether it is sufficient to study pure equilibria. If there is reason to suspect that some interesting bidder behavior can only be represented with mixed equilibria (e.g., randomized bluffing), it is necessary to be cautious about the choice of MARL algorithm. Otherwise, we recommend enforcing that policies play pure strategies by rounding them to the nearest pure strategy, deterministically playing the action with the highest probability in the policy; we refer to such rounded policies as modal. This rounding step is necessary for finding pure strategy equilibria with MARL algorithms that play actions with probabilities that approach zero, but do so exceedingly slowly, potentially never reaching zero. One such example is PPO, which rewards entropy in its loss function.
@@ 2.3. Validating and Interpreting Policies
It is important to first understand whether a MARL algorithm has converged to an equilibrium: if not, it is difficult to know whether trends in bidders’ strategies are simply caused by poor convergence. The ideal measure of convergence is the NashConv of the policy, which is defined as the sum of each agent’s regret—the utility gain they would achieve by playing a best-response, holding their opponents’ strategies fixed. A NashConv of zero thus corresponds to a Nash Equilibrium, and a positive NashConv corresponds to an $\epsilon$-Nash equilibrium, where $\epsilon$ is no greater than the NashConv. Unfortunately, NashConv is nontrivial to compute, as it requires requires finding the optimal solution to a single-agent RL problem for each player, a task that requires an exhaustive search. Thus, while it would be ideal to checkpoint the trained policy periodically and to analyze only the checkpoint with the smallest NashConv (which need not be the latest checkpoint), this is infeasible on all but the smallest of games.
When it is not possible to compute NashConv exactly, it can still be valuable to approximate it by computing approximate best responses. One way is to restrict the action space of the game to a smaller set of heuristics (such as in Section 2.1.1) and compute the best response on this restricted game. Another is to run a traditional single-agent RL algorithm, which has no guarantee of finding the optimal policy, but may nonetheless find a good policy. These approximate best responses give a lower bound on a player’s regret: if one strategy would have increased their utility, then their best response must too. While this approach cannot guarantee that a strategy profile is an equilibrium, it can still be useful to have some evidence that each player’s strategy is not easy to improve upon.
@@
We have already stated some refinements above that rule out certain equilibria. Then, if a single equilibrium remains, one can simply proceed to study its strategy profile or check how its outcomes are affected by changes to the auction. However, a likely outcome is to wind up with a set of equilibria, making the analysis murkier: should an analyst treat one equilibrium as more important or plausible than another?
Without any other desiderata, we believe it is difficult to rank equilibria by their plausibility or frequency, as the exact choice of MARL algorithm biases which equilibria are likely to be found in ways that are difficult to predict. When faced with multiple equilibria, we advocate that no equilibrium is inherently more plausible than others. Thus, we recommend looking at the range of possible strategies or outcomes represented by the set of equilibria, rather than looking at e.g., the mean outcome across all samples, or only studying a single sample.
@@ 3.1. Clock Auctions
Clock auctions are an iterative combinatorial auction format that has been used to allocate spectrum, e.g., in the FCC’s Auction 102 (FCC, 2019) and the Canadian 3500 MHz auction (ISED, 2021). Here, we describe a simplified auction format that draws heavily upon these real spectrum auctions; we provide a formal description of the auction in Appendix B.
In a clock auction for allocating spectrum, an auctioneer sells indivisible radio licenses to a number of bidders, with each bidder representing a telecom company. The licenses are divided into products, with each product corresponding to an equivalence class of geographic region and spectrum quality; there are typically several units of supply available for each product. The auctioneer sets an opening price for each product. Then, the auction proceeds over a series of rounds. In each round, bidders submit a vector-valued bid representing the number of units of each product they want to buy at the current prices. If any product is overdemanded—that is, has more aggregate demand than supply—then the auctioneer increases the price on each overdemanded product, informs all bidders of every product’s aggregate demand and new price, and proceeds to the next round of the auction. Otherwise, the auction ends, with each bidder winning the units they bid for, paying their final prices.
@@
We now use our methodology to evaluate a potential design choice: when multiple bidders simultaneously request to drop licenses, how should the auctioneer determine which bids to process first? Processing every bidder’s drop request could change a product from being over- to under-demanded, which the auction rules disallow. To avoid leaving licenses unsold, the auctioneer must instead process the drop requests sequentially. In practice (e.g., in the Canadian 3500 MHz auction), these ties are handled by randomly ordering the queue of requests and processing entire drop requests in this order, letting bidders drop as many licenses as they requested, until no excess demand remains. This “drop-by-bidder” mechanism can have high variance, processing either all or none of each bidder’s submitted bid. We compare this mechanism to an alternative where a request to drop several licenses of a product is represented in the queue as multiple requests to drop an individual license. This alternative “drop-by-license” mechanism reduces variance, making it most likely that each bidder will have some of their drop requests processed. These bid processing mechanisms are well-suited for our computational methodology because they require a large amount of case-based reasoning to study, making them overwhelming to analyze with traditional methods but ideal for reinforcement learning.
@@ 3.2. Defining the Environment
We consider auctions with two bidders and a single geographic region with two products: an unencumbered [^3] In spectrum auctions, certain licenses are encumbered, only covering a fraction of the MHz-Pop that a full, unencumbered license would. Encumbered licenses are usually priced at a discount to reflect their poorer coverage. product with one license of supply, and an encumbered product with four licenses, each of which only gives 60% of the bandwidth of an unencumbered license. Each license has eligibility points and an opening price proportional to the amount of bandwidth available. We focus on a specific state midway through the auction [^4] We do not assume this particular midway point is part of any equilibrium of the game starting from the first round. We are interested in equilibrium behavior starting from this point.: we assume that each bidder has previously bid for three of the encumbered licenses for several rounds, making it more expensive per MHz of spectrum than the unencumbered license. The bidders have a strategic choice to make: continue bidding on the more expensive product, or attempt to switch to the cheaper one (e.g., dropping two encumbered licenses to pick up one unencumbered license). The bid processing mechanism is used if both bidders attempt to switch. In the drop-by-bidder mechanism, the auctioneer chooses a random order over the bidders and processes all of one bidders’ drops in order, allowing one random bidder to switch while the other’s bid is unprocessed. In the drop-by-license mechanism, they instead process the individual requested drops in a random order, most often ending with both bidders dropping one encumbered license; in this case, the activity rule prevents both bidders from picking up the unencumbered license. [^5] Real auctions often include a grace period rule, giving bidders a chance to regain lost activity in the following round. Our model does not include a grace period. However, notice that in this strategic setting, if both players lose activity, the auction is guaranteed to end, and a grace period rule would have no effect in these cases.
We set the opening start-of-round prices to \$12 million on the unencumbered product and \$7 million on the encumbered product. We set the clock speed—the price increase after each round—to 5% of the current price. Thus, after two rounds of bidding for the encumbered product, its price increased by two clock increments to \$7.72 million.
@@
We assume that each bidder has a finite number of types, with each type being equally likely. Each type has a quasilinear utility function, meaning that they have a value for each bundle, and their utility for winning a bundle is its value minus its price paid. We choose these values by drawing them from the Spectrum Auction Test Suite (SATS) value model (Weiss et al., 2017), informed by the 2014 Canadian spectrum auction. Within a region, SATS models bidders as having a critical amount of spectrum which they need in order to provide a high quality of service, and low marginal value when they have much more or less than this critical amount. Bidders are thus described by two parameters: their value per subscriber, the amount of value they would gain from winning all of the spectrum, and their market share, the fraction of spectrum which they need to obtain half of this maximum value. Their values are then sigmoidal with respect to the fraction of spectrum they win, with high marginal values within $\pm$15of their market share, and low marginal values outside of this range. [^6] We note a similar sigmoidal modeling choice in (Bichler et al., 2022), which studies the FCC’s C-band. For each type, we draw their value per subscriber uniformly at random between \$20 million and \$30 million, and their market share uniformly at random between 35% and 50%. Note that these non-linear value functions can make “switch” bids risky, as bidders can lose a large amount of value if a switch bid is only partially processed.
In preliminary experiments, we found that some samples from this value model led to games that were either strategically uninteresting (e.g., had a bidder too weak to have a chance of winning a single license) or had game trees too large to be feasible (e.g., could last 20 rounds or more). To guard against these problems, we ran rejection sampling, resampling values that would exhibit one of these problems. We generated values for bidders having 1, 2, 3, 5, and 7 types, generating 5 value profiles for each number of types. We provide the precise details of the rejection sampling process and an example of bidders’ value functions in Appendix C.
@@
With no further changes, these auctions can admit many different actions that lead to the same terminal rewards: for example, attempting to drop from a demand of 3 to a demand of 0 often has the same effect as dropping to a demand of 1. We address this problem by adding secondary rewards (as described in Section 2.2.2). Specifically, each round, we subtract a penalty between zero and one from each player’s utility. A penalty of zero corresponds to choosing the most profitable bundle at the current prices, and a penalty of one to the least profitable; other bundles’ penalties interpolate between these two extremes. These secondary rewards can be understood as giving bidders an incentive to prefer straightforward bids and shorter auctions in the absence of any other differences, breaking equalities between actions’ terminal rewards. These penalties are removed after training, and all of our reported NashConv values are with respect to the unmodified game.

### v2

@@ 2.2. Finding Equilibria
Multi-agent reinforcement learning is a broad field with a wide landscape of methods. Here, we describe just two key dimensions along which MARL algorithms vary. For a comprehensive overview of the field, see, e.g., Albrecht et al. (2024).
The first dimension is whether agents’ policies have a tabular or function approximation representation. In tabular methods, each information state in the game is represented separately; bidders faced with a new information state do not reason about its similarity to other states with which they have previous experience, instead learning afresh. This is a double-edged sword: it does not constrain players’ behavior in any particular way, allowing them to learn flexible strategies, but causes erratic behavior in rarely-encountered states. With function approximation (i.e., deep reinforcement learning), each information state is represented by a feature vector and the policy is represented by a parametric function, such as a deep neural network. This allows agents to generalize from one part of the game tree to another, but can be more difficult to train reliably.
A second key dimension distinguishing MARL algorithms is whether they explore counterfactual actions. Many algorithms from single-agent reinforcement learning, such as policy-gradient methods (Sutton et al., 1999), sample a single path through the game tree at each iteration, only learning about their rewards along that path. While it is possible to use these algorithms in multi-agent settings by running them independently for each player, they can struggle due to focusing most of their training effort on a small fraction of the game tree. At the other extreme, algorithms such as Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007) enumerate every information state in each iteration. This extra coverage allows these algorithms to compare rewards across different actions, but their runtime quickly becomes infeasible for larger games. Algorithms can fit between these two extremes: for example, external-sampling Monte-Carlo CFR (MCCFR) (Lanctot et al., 2009) enumerates one player’s actions but samples opponent actions and chance outcomes.
For large games, it can be infeasible to store a tabular policy. However, for smaller games, we see little reason not to use a tabular method that explores counterfactual actions, such as CFR; at the very least, we recommend such algorithms as a baseline for comparing to other approaches.
@@
Complex auctions typically admit multiple equilibria. When using non-deterministic MARL algorithms, a simple way to find multiple equilibria is to run the algorithm with multiple random seeds, which influence both its random initialization and any random events within the environment, causing it to take a different learning trajectory on each run. We thus recommend non-deterministic algorithms, e.g., preferring Monte-Carlo sampling variants of CFR over the vanilla algorithm. We speculate about potential approaches to identifying multiple equilibria with deterministic algorithms in Section 5.2.2.
When two or more actions lead to the same terminal reward, RL algorithms have no reason to prefer one action over another, and equilibria multiply. This can be detrimental to both convergence and auction outcomes. For example, consider a bidder facing an incredibly strong opponent that will price them out of the auction no matter how they behave in early rounds. Since they cannot impact their own utility, they can play any strategy in equilibrium, ranging from immediately dropping out of the auction to raising their opponent’s prices up to their own value. Their arbitrary choice will impact many auction outcomes of interest (e.g., auction length, revenue). We therefore advocate for adding small, “secondary” rewards to the RL environment, biasing agents toward certain strategies (e.g., preferring to drop out than to stay in) when all else is equal.
@@
An analyst must decide whether to focus solely on pure equilibria or to also consider mixed equilibria. There are several advantages to studying pure equilibria. First, working with pure equilibria is computationally beneficial: restricting to pure strategies greatly reduces the strategy space, tending to make it easier to train, validate, and evaluate policies. Second, some MARL algorithms may not be able to converge to mixed equilibria at all. For a broad class of MARL algorithms, the only stable points of the learning dynamics are pure equilibria (Flokas et al., 2020). While there is no guarantee that a game will have any pure equilibria, in our experiments, we were able to find a pure equilibrium in the vast majority of our games.
We recommend considering whether it is sufficient to study pure equilibria. If it is critical to allow bidders to play mixed strategies (e.g., to allow randomized bluffing), one must be cautious about the choice of MARL algorithm. Otherwise, given that the outcome of a MARL algorithm is typically mixed, we recommend rounding learned policies to the nearest pure strategy, deterministically playing the action with the highest probability; we refer to such rounded policies as modal. This rounding step is important for MARL algorithms that play actions with probabilities that approach zero, but do so slowly, potentially never reaching zero. One such example is PPO, which rewards entropy in its loss function.
@@ 2.3. Validating and Interpreting Policies
It is important to understand whether a MARL algorithm has converged to an equilibrium: if not, it is difficult to know whether trends in bidders’ strategies are simply caused by poor convergence. The ideal measure of convergence is the NashConv of the policy, which is defined as the sum of each agent’s regret—the utility gain they would achieve by playing a best-response, holding their opponents’ strategies fixed. A NashConv of zero thus corresponds to a Nash Equilibrium, and a positive NashConv corresponds to an $\epsilon$-Nash equilibrium, where $\epsilon$ is no greater than the NashConv. Unfortunately, NashConv is nontrivial to compute, as it requires requires finding the optimal solution to a single-agent RL problem for each player, a task that requires an exhaustive search. Thus, while it would be ideal to checkpoint the trained policy periodically and to analyze only the checkpoint with the smallest NashConv (which need not be the latest checkpoint), this is infeasible on all but the smallest of games.
When it is not possible to compute NashConv exactly, it can still be valuable to estimate it by computing approximate best responses. One way to do so is to restrict the action space of the game to a smaller set of heuristics (such as in Section 2.1.1) and compute best responses on this restricted game. Another is to run a traditional single-agent RL algorithm, which may find a good policy even if it has no guarantee of finding the optimal policy. These approximate best responses give a lower bound on a player’s regret, as the true best response must increase utility by at least as much as the found strategy. While this approach cannot guarantee that a strategy profile is an equilibrium (because it can fail to find optimal best responses), it can still be useful to have some evidence that each player’s strategy is not easy to improve upon.
@@
We have already stated some refinements above that rule out certain equilibria. If a single equilibrium remains for each game, one can simply study its strategy profile or check how its outcomes are affected by changes to the auction. However, the analysis is murkier if multiple equilibria remain: should an analyst treat one equilibrium as more important than another? Should the frequency with which a given equilibrium is observed have any bearing on its importance? With no other desiderata claiming that one equilibrium is more plausible than another, we believe an analyst should not rank equilibria by their frequency, as the exact choice of MARL algorithm biases which equilibria are likely to be found. Thus, we recommend looking at the range of possible strategies or outcomes represented by the set of equilibria, rather than looking at e.g., the mean outcome across all samples, or only studying a single sample.
@@ 3.1. Clock Auctions
Clock auctions are an iterative combinatorial auction format that have been used to allocate spectrum, e.g., in the FCC’s Auction 102 (FCC, 2019) and the Canadian 3500 MHz auction (ISED, 2021). Here, we describe a simplified auction format that draws heavily upon these real spectrum auctions; we provide a formal description in Appendix B.
In a clock auction for allocating spectrum, an auctioneer sells indivisible radio licenses to a number of bidders. The licenses are divided into products, each corresponding to an equivalence class of geography and spectrum quality; there are typically several units of supply available for each product. The auctioneer sets an opening price for each product. Then, the auction proceeds over a series of rounds. In each round, bidders submit a vector-valued bid representing the number of units of each product they want to buy at the current prices. If any product is overdemanded—that is, has more aggregate demand than supply—then the auctioneer increases the price on each overdemanded product, informs all bidders of every product’s aggregate demand and new price, and proceeds to the next round of the auction. Otherwise, the auction ends, with each bidder winning the units they bid for, paying their final prices.
@@
We now use our methodology to evaluate a potential design choice: when multiple bidders simultaneously request to drop licenses, how should the auctioneer determine which bids to process first? Processing every bidder’s drop request could change a product from being over- to under-demanded, which the auction rules disallow. To avoid leaving licenses unsold, the auctioneer must instead process the drop requests sequentially. In practice (e.g., in the Canadian 3500 MHz auction), these ties are handled by randomly ordering the queue of requests and processing entire drop requests in this order, letting bidders drop as many licenses as they requested, until no excess demand remains. This “drop-by-bidder” mechanism can have high variance, processing either all or none of each bidder’s submitted bid. We compare this mechanism to an alternative where a request to drop several licenses of a product is represented in the queue as multiple requests to drop an individual license. This alternative “drop-by-license” mechanism reduces variance, making it possible for each bidder to have some of their drop requests processed. These bid processing mechanisms are well-suited for our computational methodology because they require a large amount of case-based reasoning to study, making them overwhelming to analyze with traditional methods but ideal for reinforcement learning.
@@ 3.2. Defining the Environment
We consider auctions with two bidders and two products: an unencumbered [^3] An encumbered license only covers a fraction of the bandwidth that a full, unencumbered license would. Encumbered licenses are usually priced at a discount to reflect their poorer coverage. product with one license of supply, and an encumbered product with four licenses, each of which only gives 60% of the bandwidth of an unencumbered license. Each license has eligibility points and an opening price proportional to its bandwidth. We focus on a specific state midway through the auction [^4] We do not assume this particular midway point is part of any equilibrium of the game starting from the first round. We are interested in equilibrium behavior starting from this point.: we assume that each bidder has previously bid for three of the encumbered licenses for several rounds, making it more expensive per MHz of spectrum than the unencumbered license. The bidders have a strategic choice to make: continue bidding on the more expensive product, or attempt to switch to the cheaper one (e.g., dropping two encumbered licenses to pick up one unencumbered license). The bid processing mechanism is triggered if both bidders attempt to switch. In the drop-by-bidder mechanism, the auctioneer chooses a random order over the bidders and processes all of one bidders’ drops in order, allowing one bidder to switch while the other’s bid is unprocessed. In the drop-by-license mechanism, the auctioneer instead processes the individual requested drops in a random order, most often ending with both bidders dropping one encumbered license. In this case, the activity rule prevents both bidders from picking up the unencumbered license. [^5] Real auctions often include a grace period rule, giving bidders a chance to regain lost activity in the following round. We do not model a grace period. However, notice that in this strategic setting, if both players lose activity, the auction is guaranteed to end, and a grace period rule would have no effect in these cases.
We set the opening start-of-round prices to \$12 million on the unencumbered product and \$7 million on the encumbered product. We set the clock speed—the price increase after each round—to 5% of the current price. Thus, after two rounds of bidding for the encumbered product, its price is \$7.72 million.
@@
We assume that each bidder has a finite number of equally likely types. Each type has a quasilinear utility function, meaning their utility for winning a bundle is its value minus its price paid. We choose these values by drawing them from the MRVM value model (Weiss et al., 2017), informed by the 2014 Canadian spectrum auction. MRVM models bidders as having a critical amount of spectrum which they need in order to provide a high quality of service, and low marginal value when they have much more or less than this critical amount. Bidders are thus described by two parameters: their value per subscriber, the amount of value they would gain from winning all of the spectrum, and their market share, the fraction of spectrum which they need to obtain half of this maximum value. Their values are then sigmoidal with respect to the fraction of spectrum they win, with high marginal values within $\pm$15of their market share, and low marginal values outside of this range. [^6] We note a similar sigmoidal modeling choice in (Bichler et al., 2022), which studies the FCC’s C-band. For each type, we draw their value per subscriber uniformly at random between \$20 million and \$30 million, and their market share uniformly at random between 35% and 50%. Note that these non-linear value functions can make “switch” bids risky, as bidders can lose a large amount of value if a switch bid is only partially processed.
In preliminary experiments, we found that some samples from this value model led to games that were either strategically uninteresting (e.g., had a bidder too weak to have a chance of winning a single license) or had game trees too large to be feasible (e.g., could last 20 rounds or more). To guard against these problems, we ran rejection sampling, resampling values that exhibited one of these problems. We generated values for bidders having 1, 2, 3, 5, and 7 types, generating 5 value profiles for each number of types. We provide the precise details of the rejection sampling process and an example of bidders’ value functions in Appendix C.
@@
With no further changes, different actions can lead to the exact same terminal rewards: for example, when a product’s demand is already at supply, attempting to drop one license often has the same effect as dropping all of the licenses. We address this problem by adding secondary rewards (as described in Section 2.2.2). Specifically, each round, we subtract a penalty between zero and one from each player’s utility. A penalty of zero corresponds to choosing the most profitable bundle at the current prices, and a penalty of one to the least profitable; other bundles interpolate between these two extremes. These secondary rewards can be understood as giving bidders an incentive to prefer straightforward bids and shorter auctions all else being equal. These penalties are removed after training, and all of our reported NashConv values are with respect to the unmodified game.